\begin{thebibliography}{30}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Abbes et~al.(2020)Abbes, Zaghouani, El-Hardlo, and
  Ashour}]{abbes-etal-2020-daict}
Ines Abbes, Wajdi Zaghouani, Omaima El-Hardlo, and Faten Ashour. 2020.
\newblock \href {https://www.aclweb.org/anthology/2020.lrec-1.768} {{DAICT}: A
  dialectal {A}rabic irony corpus extracted from {T}witter}.
\newblock In \emph{Proceedings of the 12th Language Resources and Evaluation
  Conference}, pages 6265--6271, Marseille, France. European Language Resources
  Association.

\bibitem[{Abdelali et~al.(2016)Abdelali, Darwish, Durrani, and
  Mubarak}]{abdelali-etal-2016-farasa}
Ahmed Abdelali, Kareem Darwish, Nadir Durrani, and Hamdy Mubarak. 2016.
\newblock \href {https://doi.org/10.18653/v1/N16-3003} {{F}arasa: A fast and
  furious segmenter for {A}rabic}.
\newblock In \emph{Proceedings of the 2016 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Demonstrations},
  pages 11--16, San Diego, California. Association for Computational
  Linguistics.

\bibitem[{Abdul-Mageed et~al.(2011)Abdul-Mageed, Diab, and
  Korayem}]{abdul-mageed-etal-2011-subjectivity}
Muhammad Abdul-Mageed, Mona Diab, and Mohammed Korayem. 2011.
\newblock \href {https://www.aclweb.org/anthology/P11-2103} {Subjectivity and
  sentiment analysis of {M}odern {S}tandard {A}rabic}.
\newblock In \emph{Proceedings of the 49th Annual Meeting of the Association
  for Computational Linguistics: Human Language Technologies}, pages 587--591,
  Portland, Oregon, USA. Association for Computational Linguistics.

\bibitem[{Abercrombie and Hovy(2016)}]{abercrombie-hovy-2016-putting}
Gavin Abercrombie and Dirk Hovy. 2016.
\newblock \href {https://doi.org/10.18653/v1/P16-3016} {Putting sarcasm
  detection into context: The effects of class imbalance and manual labelling
  on supervised machine classification of {T}witter conversations}.
\newblock In \emph{Proceedings of the {ACL} 2016 Student Research Workshop},
  pages 107--113, Berlin, Germany. Association for Computational Linguistics.

\bibitem[{Abu~Farha and Magdy(2020)}]{abu-farha-magdy-2020-arabic}
Ibrahim Abu~Farha and Walid Magdy. 2020.
\newblock \href {https://www.aclweb.org/anthology/2020.osact-1.5} {From
  {A}rabic sentiment analysis to sarcasm detection: The {A}r{S}arcasm dataset}.
\newblock In \emph{Proceedings of the 4th Workshop on Open-Source Arabic
  Corpora and Processing Tools, with a Shared Task on Offensive Language
  Detection}, pages 32--39, Marseille, France. European Language Resource
  Association.

\bibitem[{Abu~Farha et~al.(2021)Abu~Farha, Zaghouani, and
  Magdy}]{abufarha-etal-2021-arsarcasm-v2}
Ibrahim Abu~Farha, Wajdi Zaghouani, and Walid Magdy. 2021.
\newblock Overview of the wanlp 2021 shared task on sarcasm and sentiment
  detection in arabic.
\newblock In \emph{Proceedings of the Sixth Arabic Natural Language Processing
  Workshop}.

\bibitem[{Amir et~al.(2016)Amir, Wallace, Lyu, Carvalho, and
  Silva}]{amir-etal-2016-modelling}
Silvio Amir, Byron~C. Wallace, Hao Lyu, Paula Carvalho, and M{\'a}rio~J. Silva.
  2016.
\newblock \href {https://doi.org/10.18653/v1/K16-1017} {Modelling context with
  user embeddings for sarcasm detection in social media}.
\newblock In \emph{Proceedings of The 20th {SIGNLL} Conference on Computational
  Natural Language Learning}, pages 167--177, Berlin, Germany. Association for
  Computational Linguistics.

\bibitem[{Antoun et~al.()Antoun, Baly, and Hajj}]{antoun2020arabert}
Wissam Antoun, Fady Baly, and Hazem Hajj.
\newblock Arabert: Transformer-based model for arabic language understanding.
\newblock In \emph{LREC 2020 Workshop Language Resources and Evaluation
  Conference 11--16 May 2020}, page~9.

\bibitem[{Antoun et~al.(2020)Antoun, Baly, and Hajj}]{antoun2020araelectra}
Wissam Antoun, Fady Baly, and Hazem Hajj. 2020.
\newblock \href {http://arxiv.org/abs/2012.15516} {Araelectra: Pre-training
  text discriminators for arabic language understanding}.

\bibitem[{Barbieri et~al.(2014{\natexlab{a}})Barbieri, Ronzano, and
  Saggion}]{barbieri}
Francesco Barbieri, Francesco Ronzano, and Horacio Saggion. 2014{\natexlab{a}}.
\newblock Italian irony detection in twitter: a first approach.

\bibitem[{Barbieri et~al.(2014{\natexlab{b}})Barbieri, Saggion, and
  Ronzano}]{barbieri-etal-2014-modelling}
Francesco Barbieri, Horacio Saggion, and Francesco Ronzano. 2014{\natexlab{b}}.
\newblock \href {https://doi.org/10.3115/v1/W14-2609} {Modelling sarcasm in
  {T}witter, a novel approach}.
\newblock In \emph{Proceedings of the 5th Workshop on Computational Approaches
  to Subjectivity, Sentiment and Social Media Analysis}, pages 50--58,
  Baltimore, Maryland. Association for Computational Linguistics.

\bibitem[{Bengio et~al.(2003)Bengio, Ducharme, Vincent, and Janvin}]{bengio}
Yoshua Bengio, R\'{e}jean Ducharme, Pascal Vincent, and Christian Janvin. 2003.
\newblock A neural probabilistic language model.
\newblock \emph{J. Mach. Learn. Res.}, 3(null):1137–1155.

\bibitem[{Chang et~al.(2020)Chang, Yu, Zhong, Yang, and
  Dhillon}]{chang2020taming}
Wei-Cheng Chang, Hsiang-Fu Yu, Kai Zhong, Yiming Yang, and Inderjit Dhillon.
  2020.
\newblock \href {http://arxiv.org/abs/1905.02331} {Taming pretrained
  transformers for extreme multi-label text classification}.

\bibitem[{Darwish(2014)}]{darwish}
Kareem Darwish. 2014.
\newblock \href {https://doi.org/10.1561/1500000031} {Arabic information
  retrieval}.
\newblock \emph{Foundations and Trends® in Information Retrieval}, 7:239--342.

\bibitem[{Filatova(2012)}]{filatova-2012-irony}
Elena Filatova. 2012.
\newblock \href
  {http://www.lrec-conf.org/proceedings/lrec2012/pdf/661_Paper.pdf} {Irony and
  sarcasm: Corpus generation and analysis using crowdsourcing}.
\newblock In \emph{Proceedings of the Eighth International Conference on
  Language Resources and Evaluation ({LREC}'12)}, pages 392--398, Istanbul,
  Turkey. European Language Resources Association (ELRA).

\bibitem[{Ghanem et~al.(2019)Ghanem, Karoui, Benamara, Moriceau, and
  Rosso}]{10.1145/3368567.3368585}
Bilal Ghanem, Jihen Karoui, Farah Benamara, V\'{e}ronique Moriceau, and Paolo
  Rosso. 2019.
\newblock \href {https://doi.org/10.1145/3368567.3368585} {Idat at fire2019:
  Overview of the track on irony detection in arabic tweets}.
\newblock In \emph{Proceedings of the 11th Forum for Information Retrieval
  Evaluation}, FIRE '19, page 10–13, New York, NY, USA. Association for
  Computing Machinery.

\bibitem[{Ghosh et~al.(2015)Ghosh, Guo, and
  Muresan}]{ghosh-etal-2015-sarcastic}
Debanjan Ghosh, Weiwei Guo, and Smaranda Muresan. 2015.
\newblock \href {https://doi.org/10.18653/v1/D15-1116} {Sarcastic or not: Word
  embeddings to predict the literal or sarcastic meaning of words}.
\newblock In \emph{Proceedings of the 2015 Conference on Empirical Methods in
  Natural Language Processing}, pages 1003--1012, Lisbon, Portugal. Association
  for Computational Linguistics.

\bibitem[{Graves et~al.(2013)Graves, rahman Mohamed, and Hinton}]{graves}
Alex Graves, Abdel rahman Mohamed, and Geoffrey Hinton. 2013.
\newblock \href {http://arxiv.org/abs/1303.5778} {Speech recognition with deep
  recurrent neural networks}.

\bibitem[{Hussein(2018)}]{HUSSEIN2018330}
Doaa Mohey El-Din~Mohamed Hussein. 2018.
\newblock \href {https://doi.org/https://doi.org/10.1016/j.jksues.2016.04.002}
  {A survey on sentiment analysis challenges}.
\newblock \emph{Journal of King Saud University - Engineering Sciences},
  30(4):330--338.

\bibitem[{Joshi et~al.(2017)Joshi, Bhattacharyya, and Carman}]{10.1145/3124420}
Aditya Joshi, Pushpak Bhattacharyya, and Mark~J. Carman. 2017.
\newblock \href {https://doi.org/10.1145/3124420} {Automatic sarcasm detection:
  A survey}.
\newblock \emph{ACM Comput. Surv.}, 50(5).

\bibitem[{Joshi et~al.(2015)Joshi, Sharma, and
  Bhattacharyya}]{joshi-etal-2015-harnessing}
Aditya Joshi, Vinita Sharma, and Pushpak Bhattacharyya. 2015.
\newblock \href {https://doi.org/10.3115/v1/P15-2124} {Harnessing context
  incongruity for sarcasm detection}.
\newblock In \emph{Proceedings of the 53rd Annual Meeting of the Association
  for Computational Linguistics and the 7th International Joint Conference on
  Natural Language Processing (Volume 2: Short Papers)}, pages 757--762,
  Beijing, China. Association for Computational Linguistics.

\bibitem[{Joshi et~al.(2016)Joshi, Tripathi, Bhattacharyya, and
  Carman}]{joshi-etal-2016-harnessing}
Aditya Joshi, Vaibhav Tripathi, Pushpak Bhattacharyya, and Mark~J. Carman.
  2016.
\newblock \href {https://doi.org/10.18653/v1/K16-1015} {Harnessing sequence
  labeling for sarcasm detection in dialogue from {TV} series {`}{F}riends{'}}.
\newblock In \emph{Proceedings of The 20th {SIGNLL} Conference on Computational
  Natural Language Learning}, pages 146--155, Berlin, Germany. Association for
  Computational Linguistics.

\bibitem[{Karoui et~al.(2017)Karoui, Zitoune, and Moriceau}]{KAROUI2017161}
Jihen Karoui, Farah~Banamara Zitoune, and Véronique Moriceau. 2017.
\newblock \href {https://doi.org/https://doi.org/10.1016/j.procs.2017.10.105}
  {Soukhria: Towards an irony detection system for arabic in social media}.
\newblock \emph{Procedia Computer Science}, 117:161--168.
\newblock Arabic Computational Linguistics.

\bibitem[{Krizhevsky et~al.(2012)Krizhevsky, Sutskever, and
  Hinton}]{krizhevsky}
Alex Krizhevsky, Ilya Sutskever, and Geoffrey~E. Hinton. 2012.
\newblock Imagenet classification with deep convolutional neural networks.
\newblock In \emph{Proceedings of the 25th International Conference on Neural
  Information Processing Systems - Volume 1}, NIPS'12, page 1097–1105, Red
  Hook, NY, USA. Curran Associates Inc.

\bibitem[{Liu(2012)}]{liu}
Bing Liu. 2012.
\newblock \href {https://doi.org/10.2200/S00416ED1V01Y201204HLT016} {Sentiment
  analysis and opinion mining}.
\newblock volume~5.

\bibitem[{Mikolov et~al.(2013)Mikolov, Sutskever, Chen, Corrado, and
  Dean}]{mikolov}
Tomas Mikolov, Ilya Sutskever, Kai Chen, Greg Corrado, and Jeffrey Dean. 2013.
\newblock Distributed representations of words and phrases and their
  compositionality.
\newblock In \emph{Proceedings of the 26th International Conference on Neural
  Information Processing Systems - Volume 2}, NIPS'13, page 3111–3119, Red
  Hook, NY, USA. Curran Associates Inc.

\bibitem[{Rajadesingan et~al.(2015)Rajadesingan, Zafarani, and
  Liu}]{10.1145/2684822.2685316}
Ashwin Rajadesingan, Reza Zafarani, and Huan Liu. 2015.
\newblock \href {https://doi.org/10.1145/2684822.2685316} {Sarcasm detection on
  twitter: A behavioral modeling approach}.
\newblock In \emph{Proceedings of the Eighth ACM International Conference on
  Web Search and Data Mining}, WSDM '15, page 97–106, New York, NY, USA.
  Association for Computing Machinery.

\bibitem[{Shaalan(2010)}]{10.1007/s10590-011-9087-8}
Khaled Shaalan. 2010.
\newblock \href {https://doi.org/10.1007/s10590-011-9087-8} {Nizar y. habash,
  introduction to arabic natural language processing (synthesis lectures on
  human language technologies)}.
\newblock \emph{Machine Translation}, 24(3–4):285–289.

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones,
  Gomez, Kaiser, and Polosukhin}]{vaswani2017attention}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N. Gomez, Lukasz Kaiser, and Illia Polosukhin. 2017.
\newblock \href {http://arxiv.org/abs/1706.03762} {Attention is all you need}.

\bibitem[{Yih et~al.(2011)Yih, Toutanova, Platt, and Meek}]{yih2011}
Wen-tau Yih, Kristina Toutanova, John~C. Platt, and Christopher Meek. 2011.
\newblock \href {https://www.aclweb.org/anthology/W11-0329} {Learning
  discriminative projections for text similarity measures}.
\newblock In \emph{Proceedings of the Fifteenth Conference on Computational
  Natural Language Learning}, pages 247--256, Portland, Oregon, USA.
  Association for Computational Linguistics.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt,a4paper]{article}
\usepackage{arabtex}
\usepackage{utf8}
\usepackage[hyperref]{eacl2021}
\usepackage{times}
\usepackage{multirow}
\usepackage{latexsym}

\usepackage{microtype}

\aclfinalcopy 

\title{AraBERT and Farasa Segmentation Based Approach For Sarcasm and Sentiment Detection in Arabic Tweets}

\author{Anshul Wadhawan \\
  Flipkart Private Limited \\
  \texttt{anshul.wadhwan@flipkart.com} \\} 

\date{}

\begin{document}
\setcode{utf8}
\maketitle

\begin{abstract}

This paper presents our strategy to tackle the EACL WANLP-2021 Shared Task 2: Sarcasm and Sentiment Detection. One of the subtasks aims at developing a system that identifies whether a given Arabic tweet is sarcastic in nature or not, while the other aims to identify the sentiment of the Arabic tweet. We approach the task in two steps. The first step involves pre processing the provided ArSarcasm-v2 dataset by performing insertions, deletions and segmentation operations on various parts of the text. The second step involves experimenting with multiple variants of two transformer based models, AraELECTRA and AraBERT. Our final approach was ranked seventh and fourth in the Sarcasm and Sentiment Detection subtasks respectively.    

\end{abstract}

\section{Introduction}

During the last two decades, work on subjective language processing has been very common in literature. The work on sentiment analysis was a major theme that was pursued during this time. Sentiment Analysis is a process, according to \cite{liu}, in which we extract out and examine the emotional stance in a particular piece of text. With the introduction of user-driven networks such as websites for social media, the work on Sentiment Analysis flourished. Most of this work was based on English, while not much attention was gathered by the Arabic language. The work on Arabic Sentiment Analysis was initiated by \cite{abdul-mageed-etal-2011-subjectivity}, but as compared to English it still needs development. This can be due to the many complexities pertaining to the language, including the broad range of dialects \cite{10.1007/s10590-011-9087-8,darwish} and the complicated language morphology \cite{abdul-mageed-etal-2011-subjectivity}. Some of the prominent problems while tackling the task of Sentiment Analysis are domain dependency, negation handling, lack of sarcasm and knowledge \cite{HUSSEIN2018330}. Sarcasm is described as a form of verbal irony designed to convey disdain or ridicule \cite{10.1145/3124420}. There has been a lot of work on the identification of English sarcasm, including datasets, like the works of \cite{barbieri-etal-2014-modelling,ghosh-etal-2015-sarcastic,abercrombie-hovy-2016-putting,filatova-2012-irony,joshi-etal-2016-harnessing,barbieri} and detection systems like \cite{joshi-etal-2015-harnessing,10.1145/2684822.2685316,amir-etal-2016-modelling}. Work on Arabic sarcasm is, to our knowledge, restricted to the work of \cite{KAROUI2017161}, an irony detection task \cite{10.1145/3368567.3368585}, and sarcasm datasets by \cite{abbes-etal-2020-daict,abu-farha-magdy-2020-arabic}.

This paper puts forth the approach we applied to handle the WANLP-2021 Shared Task 2. The paper is ordered in the following manner: The problem statement, along with details of the ArSarcasm-v2 dataset are presented in Section 2. The methodology that we propose as our solution is described in Section 3. The experiments which were carried out, dataset statistics, system settings and results of the experiments are provided in Section 4. The paper ends with a brief section that talks about the conclusion and future directions of our research, in section 5.

\section{Task Definition}

The WANLP-2021 Shared Task 2 \cite{abufarha-etal-2021-arsarcasm-v2} is based on a text classification problem, based on identifying sentiment and sarcasm in Arabic tweets. The provided training and test datasets have a total of 12,548 and 3,000 tweets respectively.

The shared task is divided into two subtasks:

Subtask 1 (Sarcasm Detection): The aim is to identify whether a tweet is sarcastic or not. Given a tweet, the task is to return TRUE if there is sarcasm present in the tweet and FALSE otherwise. This is a binary classification problem. Precision/Recall/F-score/Accuracy are the evaluation metrics, where F-score of the sarcastic class is the official metric of evaluation.

Subtask 2 (Sentiment Analysis): The aim is to identify the sentiment of a tweet by assigning one of three labels (Positive, Negative, Neutral). Given a tweet, the task is to return POS if the tweet has a positive sentiment, NEG if the tweet has a negative sentiment or NEU if the tweet has a neutral sentiment. This is a multiclass classification problem. Precision/Recall/F-score/Accuracy are the evaluation metrics, where macro average of the F-score of the positive and negative classes (F-PN) is the official metric of evaluation.

\section{Methodology}

This section describes the process we employed to tackle the task. The process is divided into two steps: data preprocessing and transformer based models application. The first step involves processing the provided ArSarcasm-v2 dataset to convert it into a format better processed by the models. The second step involves experimenting with different models to decide which model performs the best for the ArSarcasm-v2 dataset. Details about these steps have been provided in the following sub-sections.

\subsection{Data Pre-Processing}

The data that is used for pre training the transformer based models is processed to create a better representation of the data. Thus, for the models to perform at their best ability, the data has to be processed in the same manner in the fine tuning process also. Raw data that is fetched from social media websites is diverse due to vast differences in expressions of opinions among users from different parts of the world. ArSarcasm-v2 dataset has these variations in different forms, evident from manual analysis. Social media users often make use of slang words, and non ascii characters such as emojis. URLs, user mentions and spelling errors are prominent in many posts on social media platforms. These attributes do not qualify as discerning features for classification tasks like Sarcasm and Sentiment Detection, and contribute to noise within the dataset. Therefore, we employ different pre processing techniques to clear this noise, so that the transformer based models only receive the relevant features. These techniques are as follows:

\begin{enumerate}
\item Remove HTML line breaks and markup, unwanted characters like emoticons, repeated characters ($>$ 2) and extra spaces.

\item Perform Farasa segmentation (for select models only) \cite{abdelali-etal-2016-farasa}.

\item Insert whitespace before and after all non Arabic digits or English Digits and Alphabet and the 2 brackets, and between words and numbers or numbers and words.

\item Replace all URLs with [ \< رابط > ], emails with [~\< بريد > ], mentions with [ \< مستخدم > ].

\end{enumerate}

\subsection{Transformer Based Models}

Deep learning methods have shown promising results in different machine learning domains such as Computer Vision \cite{krizhevsky} and Speech Recognition \cite{graves}. The traditional machine learning methods have been over taken by deep learning techniques in the recent past, because of their superior performance owing to architectures inspired by the human brain. On the lines of Natural Language Processing, most deep learning techniques have been making use of word vector representations \cite{yih2011,bengio,mikolov} mainly as a way of representing the textual inputs. These techniques are further being replaced by transformer based techniques \cite{vaswani2017attention} due to significant improvements on most NLP tasks like text classification \cite{chang2020taming}, which is the task at hand. Transformer based techniques have the ability to produce efficient embeddings as an output of the pre training process, which makes them proficient language models.

\begin{table*}[]
\begin{tabular}{|l|l|l|l|l|l|l|}
\hline
\multicolumn{1}{|c|}{\multirow{2}{*}{\textbf{Model}}} & \multicolumn{2}{c|}{\textbf{Size}}                                      & \multicolumn{1}{c|}{\multirow{2}{*}{\textbf{Pre-Segmentation}}} & \multicolumn{3}{c|}{\textbf{Dataset}}                                                                                  \\ \cline{2-3} \cline{5-7} 
\multicolumn{1}{|c|}{}                                & \multicolumn{1}{c|}{\textbf{MB}} & \multicolumn{1}{c|}{\textbf{Params}} & \multicolumn{1}{c|}{}                                           & \multicolumn{1}{c|}{\textbf{\#Sentences}} & \multicolumn{1}{c|}{\textbf{Size}} & \multicolumn{1}{c|}{\textbf{\#Words}} \\ \hline

AraBERTv0.1-base                                      & 543MB                            & 136M                                 & No                                                              & 77M                                       & 23GB                               & 2.7B                                  \\ \hline
AraBERTv1-base                                        & 543MB                            & 136M                                 & Yes                                                             & 77M                                       & 23GB                               & 2.7B                                  \\ \hline
AraBERTv0.2-base                                      & 543MB                            & 136M                                 & No                                                              & 200M                                      & 77GB                               & 8.6B                                  \\ \hline
AraBERTv0.2-large                                     & 1.38G                            & 371M                                 & No                                                              & 200M                                      & 77GB                               & 8.6B                                  \\ \hline
AraBERTv2-base                                        & 543MB                            & 136M                                 & Yes                                                             & 200M                                      & 77GB                               & 8.6B                                  \\ \hline
AraBERTv2-large                                       & 1.38G                            & 371M                                 & Yes                                                             & 200M                                      & 77GB                               & 8.6B                                  \\ \hline
\end{tabular}
\caption{Architectural Attributes of Models}
\end{table*}

\subsubsection{AraBERT} \cite{antoun2020arabert} Pre trained to handle Arabic text, AraBERT is a language model that is inspired from the Google's BERT architecture. Six variants of the same model are available for esxperimentation: AraBERTv0.2-base, AraBERTv1-base, AraBERTv0.1-base, AraBERTv2-large, AraBERTv0.2-large, and AraBERTv2-base. The architectural attributes of each of these models have been highlighted in Table 1.

\subsubsection{AraELECTRA} \cite{antoun2020araelectra} With reduced computations for pre training the transformers, ELECTRA is a method aimed towards the task of self-supervised language representation learning. ELECTRA models are inspired from the two primary components of Generative Adversarial Networks: generator and discriminator. They aim at distinguishing between real input tokens and the fake ones. These models have shown convincing state-of-the art results on Arabic QA data.

For the pretraining process of all new AraBERT and AraELECTRA models, the same data is used, which has a size of 77GB. It has in total 8,655,948,860 words or 82,232,988,358 characters or 200,095,961 lines, before Farasa segmentation. For an initial pre training dataset, several websites were crawled, which are: Assafir news articles, OSCAR unshuffled and filtered, The OSIAN Corpus, Arabic Wikipedia dump from 2020/09/01, and The 1.5B words Arabic Corpus. For newer models, a fresh dataset was developed for the pre training process, which did not include the data crawled from the above websites, but had unshuffled and properly filtered OSCAR corpus in addition to the dataset used in AraBERTv1. 

\section{Experiments}

We split the provided training dataset in a 90:10 ratio to create our training and validation splits. This is followed by experimenting with eight transformer based models, for which performance on the validation split is compared. The pre trained models are fine tuned on the training split and metrics for the validation split are calculated. The final test set predictions are made from the model that performs the best on the validation split, among the eight models. This section includes the dataset distribution, system settings, results and a brief analysis of the system, for both the subtasks: sarcasm and sentiment detection. 

\subsection{Dataset}

\begin{table}[]
\centering
\begin{tabular}{|l|l|l|l|}
\hline
\multicolumn{1}{|c|}{\multirow{2}{*}{\textbf{}}} & \multicolumn{2}{c|}{\textbf{Sarcasm}}                                    & \multicolumn{1}{c|}{\multirow{2}{*}{\textbf{Total}}} \\ \cline{2-3}
\multicolumn{1}{|c|}{}                           & \multicolumn{1}{c|}{\textbf{FALSE}} & \multicolumn{1}{c|}{\textbf{TRUE}} & \multicolumn{1}{c|}{}                                \\ \hline
Train                                            & 9356                                & 1937                               & 11293                                                \\ \hline
Dev                                              & 1024                                & 231                                & 1255                                                 \\ \hline
Total                                            & 10380                               & 2168                               & 12548                                                \\ \hline
\end{tabular}
\caption{Data Distribution w.r.t Sarcasm}
\end{table}

\begin{table}[]
\centering
\begin{tabular}{|l|l|l|l|l|}
\hline
\multicolumn{1}{|c|}{\multirow{2}{*}{\textbf{}}} & \multicolumn{3}{c|}{\textbf{Sentiment}}                                                                   & \multicolumn{1}{c|}{\multirow{2}{*}{\textbf{Total}}} \\ \cline{2-4}
\multicolumn{1}{|c|}{}                           & \multicolumn{1}{c|}{\textbf{NEG}} & \multicolumn{1}{c|}{\textbf{NEU}} & \multicolumn{1}{c|}{\textbf{POS}} & \multicolumn{1}{c|}{}                                \\ \hline
Train                                            & 4139                              & 5197                              & 1957                              & 11293                                                \\ \hline
Dev                                              & 482                               & 550                               & 223                               & 1255                                                 \\ \hline
Total                                            & 4621                              & 5747                              & 2180                              & 12548                                                \\ \hline
\end{tabular}
\caption{Data Distribution w.r.t Sentiment}
\end{table}

Tables 2 and 3 show the class wise distribution of the 90:10 training-validation splits created from the provided ArSarcasm-v2 dataset, for the tasks of sarcasm detection which has class labels TRUE and FALSE and sentiment detection which has class labels NEG, NEU and POS, respectively.

\subsection{System Settings}

\begin{table}[]
\centering
\begin{tabular}{|l|l|}
\hline
\textbf{Parameter}        & \textbf{Value}  \\ \hline
Epsilon (Adam optimizer)  & 1e-8            \\ \hline
Learning Rate             & 1e-5            \\ \hline
Batch Size (for base models)                & 40               \\ \hline
Batch Size (for large models)                & 4               \\ \hline
Maximum Sequence Length   & 256             \\ \hline
\#Epochs                  & 10               \\ \hline
\end{tabular}
\caption{Parameter Values}
\end{table}

\begin{table*}[]
\centering
\begin{tabular}{|l|l|l|l|l|l|l|l|l|}
\hline
\multicolumn{1}{|c|}{\multirow{2}{*}{\textbf{Model}}} & \multicolumn{4}{c|}{\textbf{Subtask 1}}                                                                                                & \multicolumn{4}{c|}{\textbf{Subtask 2}}                                                                                                \\ \cline{2-9} 
\multicolumn{1}{|c|}{}                                & \multicolumn{1}{c|}{\textbf{P}} & \multicolumn{1}{c|}{\textbf{R}} & \multicolumn{1}{c|}{\textbf{F1}} & \multicolumn{1}{c|}{\textbf{A}} & \multicolumn{1}{c|}{\textbf{P}} & \multicolumn{1}{c|}{\textbf{R}} & \multicolumn{1}{c|}{\textbf{F1}} & \multicolumn{1}{c|}{\textbf{A}} \\ \hline
AraBERTv0.1-base                                      & 84.15                           & 84.62                           & 84.36                            & 84.62                           & 72.38                           & 72.43                           & 72.40                            & 72.43                           \\ \hline
AraBERTv0.2-base                                      & 84.31                           & 85.42                           & 84.62                            & 85.42                           & 74.90                           & 74.98                           & 74.90                            & 74.98                           \\ \hline
AraBERTv0.2-large                                     & 84.57                           & 85.90                           & 84.63                            & 85.90                           & 76.25                           & 76.41                           & 76.20                            & 76.41                           \\ \hline
AraBERTv1-base                                        & 84.73                           & 85.02                           & 84.86                            & 85.02                           & 73.27                           & 73.55                           & 73.17                            & 73.55                           \\ \hline
AraBERTv2-base                                        & 85.39                           & 85.74                           & 85.55                            & 85.74                           & 75.93                           & 75.86                           & 75.89                            & 75.86                           \\ \hline
AraBERTv2-large                                       & 85.16                           & 86.37                           & 85.07                            & 86.37                           & 74.28                           & 74.50                           & 74.31                            & 74.50                           \\ \hline
AraELECTRA-base-generator                             & 82.72                           & 83.19                           & 82.94                            & 83.19                           & 72.45                           & 72.03                           & 71.99                            & 72.03                           \\ \hline
AraELECTRA-base-discriminator                         & 84.75                           & 85.34                           & 84.99                            & 85.34                           & 74.70                           & 74.82                           & 74.57                            & 74.82                           \\ \hline
\end{tabular}
\caption{Results on Validation Set}
\end{table*}

We make use of hugging-face\footnote{\url{https://huggingface.co/transformers/}} API to fetch the pre-trained AraBERT and AraELECTRA models. The API provides the six variants of AraBERT models by the names of bert-base-arabertv02, bert-base-arabert, bert-base-arabertv2, bert-base-arabertv01, bert-large-arabertv2, and bert-large-arabertv02, and the two variants of AraELECTRA models by the names of araelectra-base-discriminator and araelectra-base-generator. We fine tune these models on the training split with hyper parameter values specified in Table 4.

\subsection{Results and Analysis}

This section provides the detailed results obtained on the created validation and provided test sets.

\subsubsection{Validation Set Results}

The experimental results in terms of weighted Precision(P), weighted F1 scores(F1), weighted Recall(R) and Accuracy(A) on the created validation split have been depicted in Table 5, for both subtasks: Sarcasm Detection (Subtask 1) and Sentiment Detection (Subtask 2).

Our observations from Table 5 are as follows:
\begin{enumerate}

\item When comparing all the models, AraELECTRA generator model has the worst performance in terms of both F1 scores and Accuracy, for both the subtasks. This is possibly due to its forte of handling GAN related tasks rather than general classification tasks.

\item When comparing all AraBERT models, one of the large models seem to perform the best in terms of accuracy for both the subtasks. This is possibly due to superior architectures and heavier models. 

\item For Subtask 1, AraBERTv2-base has the highest weighted F1-score and for Subtask 2, AraBERTv0.2-large has the highest weighted F1-score. 

\end{enumerate}

\begin{table}[]
\centering
\begin{tabular}{|c|c|c|c|c|c|}
\hline
              & \textbf{C\textsubscript{E}} & \textbf{A} & \textbf{P} & \textbf{R} & \textbf{M-F1}\\ \hline
T\textsubscript{S}\textsuperscript{1} & 58.72  & 78.30 & 72.64 & 71.47 & 72.00\\ \hline
T\textsubscript{S}\textsuperscript{2} & 72.55 & 69.83  & 65.15 & 66.23 & 65.31\\ \hline

\end{tabular}
\caption{Official Results on Test Set}
\end{table}

\subsubsection{Test Set Results}

From the above observations, we select AraBERTv2-base for Subtask 1 and AraBERTv0.2-large for Subtask 2 as final models to formulate our officially submitted predictions. Table 6 presents the final test set results, with T\textsubscript{S}\textsuperscript{1} denoting Subtask 1, T\textsubscript{S}\textsuperscript{2} denoting Subtask 2, P denoting Precision, R denoting Recall, A denoting Accuracy, M-F1 denoting Macro-F1 score, and C\textsubscript{E} denoting the criteria of evaluation for the two subtasks (F-score of the sarcastic class for T\textsubscript{S}\textsuperscript{1}, and Macro averaged F score of positive and negative classes for T\textsubscript{S}\textsuperscript{2}). 

\section{Conclusion and Future Work}

In this paper, we present our strategy to approach the EACL WANLP-2021 Shared Task 2. We tackle the task in two steps. In the first step, the ArSarcasm-v2 dataset is pre-processed by altering different parts of text. This is followed by running experiments with multiple variants of two transformer based models pre-trained on Arabic text, AraELECTRA and AraBERT. The final submissions for the tasks of Sarcasm and Sentiment Detection are based on that variant of model which performs the best. Our approach fetches a private leaderboard rank 7 and 4 in the Sarcasm and Sentiment Detection tasks respectively. As future scope, we plan to explore other features which may be relevant for this task, and inculcate ensemble learning taking into consideration both word vector based and transformer based embeddings.

\bibliography{eacl2021}
\bibliographystyle{acl_natbib}

\end{document}